OpenStudy (anonymous):

Add or subtract terms whenever possible. -3 sqrt 162 + 2 sqrt 200 - 9 sqrt 8

OpenStudy (anonymous):

200=100*2=10^2*2 and 8 = 2*2*2=2^2*2

OpenStudy (anonymous):

my answer choices are A.-3 sqrt 2 B .-187 sqrt 2 C. 187 SQRT 2 D. -25 SQRT 2

OpenStudy (anonymous):

so you have 162 as 9^2*2, and 200 as 10^2*2m and 8 as 2^2*2 so...

OpenStudy (anonymous):

9 can come out, 10 can come out, 2 can come out. so you will have -27sqrt (2)+ 20sqrt(2)-18sqrt(2), and you can add the numbers in front if the number in sqrt is identical.

OpenStudy (anonymous):

OK its -25 sqrt 2?

OpenStudy (anonymous):

@katherine.ok could you help me with another please?

OpenStudy (anonymous):

@Rowa could you help me please

OpenStudy (anonymous):

@GoldPhenoix could you help me please?

OpenStudy (anonymous):

Multiply or divide as indicated.

OpenStudy (anonymous):

@zepdrix could you help me

OpenStudy (anonymous):



zepdrix (zepdrix):

Is this what you have written? \[\large \frac{2x}{4x+2}\times \frac{14x+7}{6}\]